# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ VII Square Root Law and Approximation of the Distribution of Noise
VII-B Square Root Law and Approximation of the Distribution of Noise
@@
  • With given $10<\delta<$ 1 and $p_{n}=c\cdot n^{-\frac{1}{2}}$ where $c$ is a constant, we can increase $n$ to satisfy any small decoding error probability $\epsilon$ without violating any proper covert constraint since the divergences between $\mathbb{P}^{(n)}_{0}$ and $\mathbb{P}^{(n)}_{1}$ will be stationary and the value depends on $c$.
@@
In this work and our previous work [24], truncated Gaussian distributions are chosen as the distributions for random coding over AWGN channels. There are two reasons of the choice. First, the way of truncation is at our choice, which will be convenient to satisfy the power constraint. The corresponding output distributions are close to Gaussian distributions with zero mean and i.i.d coordinates, hence they will lead to relatively tight normal approximations of finite blocklength throughput, which has been explained in [24]. Second, as we have explained in Section VI, truncated Gaussian distributions with proper chosen variances are good approximations of Dirac measures, which make them natural candidates on random coding for covert communications over AWGN channels. From the framework, other measures as approximations of Dirac measures can also be utilized in random coding over AWGN covert channels. Nevertheless, the number of the reliable bits from the corresponding random codes may be limited. Considering the mutual information as an approximate metric of the upper bound on the number of the transmitted bits, it is reasonable that the input distribution should be close to Gaussian.
VIII Conclusion
@@
This paper presents a comprehensive analysis of covert communication over the AWGN channel, focusing on both achievability and converse bounds under finite blocklength constraints. By characterizing the first and second order asymptotics of Gaussian random coding, we establish fundamental limits on the trade-offs between reliability, covertness, and key size. Notably, the dependence of these asymptotics on the hyper-parameter $\mu$ underscores the critical role of power constraints in practical coding designs for finite $n$. Our results reveal that the required number of key bits scales as $O(\sqrt{n})$, a conclusion that remains robust even when the noise level at the legitimate receiver is marginally lower than that at the adversary. This highlights the inherent necessity of shared secrecy resources to sustain covertness in scenarios with minimal channel asymmetry. Furthermore, the converse bounds derived here align precisely with the supremum of the achievability results over $\mu$, demonstrating the tightness of our analysis and the optimality of Gaussian random coding in this regime.
From an information-geometric perspective, the truncated Gaussian distribution emerges as the natural choice for codebook generation, balancing the trade-offs between power constraints and divergence-based covertness metrics. The Square Root Law for optimal power allocation is rigorously justified through the asymptotic behavior of divergences between Gaussian distributions and the statistical indistinguishability of noise. These findings not only unify existing theoretical frameworks but also provide practical insights for designing covert communication systems with provable guarantees.
@@ Appendix A Proof of Lemma 1
$\begin{split}\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=&\int_{\bm{x}}f^{(n)}(\bm{x})\log\frac{f^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}d\bm{x}\\ =&\int_{\bm{x}}f^{(n)}(\bm{x})\log\frac{1}{\Delta}d\bm{x}\\ =&-\log\Delta.\end{split}$ (152)
@@
$\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=$ $\frac{1}{2}\underset{\mathbb{R}^{n}}{\int}|f^{(n)}(\bm{x})-g^{(n)}(\bm{x})|d\bm{x}$
@@
Appendix B Proof of Lemma 3
@@
Consider the distributions $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ as convolution operators $\bar{\mathfrak{F}}_{n}^{(n)}$ and $\mathfrak{F}_{n}^{(n)}$ on $C(\mathbb{R}^{n})$, respectively. For bounded continuous functions $f_{0}^{(n)}(\bm{y})$ and $\hat{f}_{0}^{(n)}(\bm{y})$, we have
$\mathfrak{F}_{n}^{(n)}f_{0}^{(n)}(\bm{y})={g}_{1}^{(n)}(\bm{y}),$ (154)
$\bar{\mathfrak{F}}_{n}^{(n)}f_{0}^{(n)}(\bm{y})=\bar{g}_{1}^{(n)}(\bm{y}),$ (155)
@@
$\mathfrak{F}_{n}^{(n)}\hat{f}_{0}^{(n)}(\bm{z})=\bar{f}^{(n)}_{1}(\bm{z}),$ (156)
$\bar{\mathfrak{F}}_{n}^{(n)}\hat{f}_{0}^{(n)}(\bm{z})=\hat{f}_{1}^{(n)}(\bm{z}).$ (157)
Thus, from Corollary 1 and Theorem 1.
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\bar{g}_{1}^{(n)}(\bm{y})-{g}_{1}^{(n)}(\bm{y})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (158)
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\hat{f}_{1}^{(n)}(\bm{z})-\bar{f}^{(n)}_{1}(\bm{z})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (159)
As the function $h(x)=\log x$ is continuous, we have
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\log\bar{g}_{1}^{(n)}(\bm{y})-\log g_{1}^{(n)}(\bm{y})|$
$=$ $\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\log\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (160)
@@
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\log\hat{f}_{1}^{(n)}(\bm{z})-\log\bar{f}^{(n)}_{1}(\bm{z})|$
$=$ $\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}\log\frac{\hat{f}_{1}^{(n)}(\bm{z})}{\bar{f}^{(n)}_{1}(\bm{z})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (161)
@@
$\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty,$ (162)
$\frac{\hat{f}_{1}^{(n)}(\bm{z})}{\bar{f}^{(n)}_{1}(\bm{z})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty.$ (163)
@@
Appendix C Proof of Lemma 7
@@
Appendix D Proof of Lemma 8
@@
$\bigg{|}\mathbb{D}(Q_{\bm{Z}}^{1}\|\bar{Q}^{(n)}_{0})-\mathbb{D}(Q_{\bm{Z}}^{2}\|\bar{Q}^{(n)}_{0})\bigg{|}$
$=$ $\bigg{|}\int Q_{\bm{Z}}^{1}\log\frac{Q_{\bm{Z}}^{1}}{\bar{Q}^{(n)}_{0}}d\bm{z}-\int Q_{\bm{Z}}^{2}\log\frac{Q_{\bm{Z}}^{2}}{\bar{Q}^{(n)}_{0}}d\bm{z}\bigg{|}$
$=$ $\bigg{|}\int Q_{\bm{Z}}^{1}\log Q_{\bm{Z}}^{1}d\bm{z}-\int Q_{\bm{Z}}^{2}\log Q_{\bm{Z}}^{2}d\bm{z}$
$-\int(Q_{\bm{Z}}^{1}-Q_{\bm{Z}}^{2})\log\bar{Q}^{(n)}_{0}d\bm{z}\bigg{|}$
$=$ $\bigg{|}\int\big{(}Q_{\bm{Z}}^{1}-Q_{\bm{Z}}^{2}\big{)}\log Q_{\bm{Z}}^{1}d\bm{z}$
$+\int Q_{\bm{Z}}^{2}\log Q_{\bm{Z}}^{1}d\bm{z}-\int Q_{\bm{Z}}^{2}\log Q_{\bm{Z}}^{2}d\bm{z}$
$-\int(Q_{\bm{Z}}^{1}-Q_{\bm{Z}}^{2})\log\bar{Q}^{(n)}_{0}d\bm{z}\bigg{|}$
$\leq$ $\bigg{|}\int\big{(}Q_{\bm{Z}}^{1}-Q_{\bm{Z}}^{2}\big{)}\log\frac{Q_{\bm{Z}}^{1}}{\bar{Q}^{(n)}_{0}}d\bm{z}\bigg{|}+\bigg{|}\int Q_{\bm{Z}}^{2}\log\frac{Q_{\bm{Z}}^{2}}{Q_{\bm{Z}}^{1}}d\bm{z}\bigg{|}$ (165)
We analyze each term in (D) separately. For the first term, as the difference between $Q_{\bm{Z}}^{1}$ and $Q_{\bm{Z}}^{2}$ is the $k$-th component, we have
$\bigg{|}\int\big{(}Q_{\bm{Z}}^{1}-Q_{\bm{Z}}^{2}\big{)}\log\frac{Q_{\bm{Z}}^{1}}{\bar{Q}^{(n)}_{0}}d\bm{z}\bigg{|}$
@@
$-\|\bm{x}\|^{2}\leq\|\bm{z}\|^{2}-\|\bm{z}-\bm{x}_{i}\|^{2}\leq\|\bm{x}\|^{2}$ (167)
and each $\bm{x}$ satisfies $\mu n\Psi(n)\leq\|\bm{x}\|^{2}\leq n\Psi(n)$. For the second term, let
@@
then $Q_{\bm{Z}}^{1}$ and $Q_{\bm{Z}}^{2}$ can be expressed as
$Q_{\bm{Z}}^{1}=\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2},\,\ \ \ \,Q_{\bm{Z}}^{2}=\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}.$ (170)
@@
=$\overset{(c)}{=}$ $\frac{\|\bm{x}_{k}-\bm{x}^{\prime}_{k}\|^{2}}{2M\sigma^{2}}$
@@
where (b) holds because KL divergence $\mathbb{D}(p\|q)$ is a convex function with the pair $(p,q)$ [34], (c) follows from the close form of KL divergence between two Gaussian distributions [35], and (d) follows from the fact both $\|\bm{x}_{k}\|^{2}$ and $\|\bm{x}^{\prime}_{k}\|^{2}$ are in $[\mu^{2}n\Psi(n),n\Psi(n)]$. From (D), (D) and (D), we have
$\bigg{|}\mathbb{D}(Q_{\bm{Z}}^{1}\|\bar{Q}^{(n)}_{0})-\mathbb{D}(Q_{\bm{Z}}^{2}\|\bar{Q}^{(n)}_{0})\bigg{|}\leq\frac{4n\Psi(n)}{M\sigma^{2}}.$ (172)
Appendix E Proof of Theorem 4
From the definitions of the distributions $\mathbb{Q}^{(n)}_{in}$ and $\mathbb{P}_{0}^{(n)}$, we have
$\mathbb{D}(\mathbb{Q}^{(n)}_{in}\|\mathbb{P}_{0}^{(n)})$
$=$ $-h(\mathbb{Q}^{(n)}_{in})+E_{\mathbb{Q}^{(n)}_{in}}\left[\log\frac{1}{f_{0}^{(n)}(\bm{y})}\right]$
$=$ $-h(\mathbb{Q}^{(n)}_{in})+\frac{n}{2}\log(2\pi)+E_{\mathbb{Q}^{(n)}_{in}}\left[\frac{\sum_{i=1}^{n}y_{i}^{2}}{2}\right]$
=$\overset{(a)}{=}$ $-h(\mathbb{Q}^{(n)}_{in})+\frac{n}{2}\log(2\pi)+\frac{n(\Psi_{a}+1)}{2}$
$\overset{(b)}{\geq}$ $-\frac{1}{2}\log(2\pi e)^{n}|\bm{K}+\bm{I}_{n}|+\frac{n}{2}\log(2\pi)+\frac{n(\Psi_{a}+1)}{2}$
=$\overset{(c)}{=}$ $-\frac{1}{2}\log(2\pi e)^{n}|\bm{Q}^{T}(\bm{K}+\bm{I}_{n})\bm{Q}|+\frac{n}{2}\log(2\pi)+\frac{n(\Psi_{a}+1)}{2}$
=$\overset{(d)}{=}$ $-\frac{1}{2}\log(2\pi e)^{n}|\prod_{i=1}^{n}\lambda_{i}|+\frac{n}{2}\log(2\pi)+\frac{n(\Psi_{a}+1)}{2}$
$\overset{(e)}{\geq}$ $-\frac{1}{2}\log(2\pi e(1+\Psi_{a}))^{n}+\frac{n}{2}\log(2\pi)+\frac{n(\Psi_{a}+1)}{2}$
$=$ $\,\ \,\mathbb{D}(\mathbb{Q}^{(n)}_{G}\|\mathbb{P}_{0}^{(n)}),$ (173)
where (a) follows from (131), (b) follows from zero mean multivariate Gaussian distribution is the maximizer of the entropy functional under second moment constraints (Chapter 12 in [34]), and (c) and (d) follow from the fact that the matrix $\bm{K}+\bm{I}_{n}$ is symmetric and positive definite, hence there exists some orthogonal matrix $\bm{Q}$ such that $\bm{Q}^{T}(\bm{K}+\bm{I}_{n})\bm{Q}=diag(\lambda_{1},\cdots,\lambda_{n})$, then $|\bm{K}+\bm{I}_{n}|=\prod_{i=1}^{n}\lambda_{i}$. In addition, the inequality (e) follows from that $\frac{1}{n}Tr(\bm{K}+\bm{I}_{n})=1+\Psi_{a}=\frac{1}{n}\sum_{i=1}^{n}\lambda_{i}$ and $\prod_{i=1}^{n}\lambda_{i}$ is maximized when $\lambda_{i},i=1,\cdots,n$ are equal. Note that the right hand side of the inequality of (b) is $\mathbb{D}(\hat{\mathbb{Q}}^{(n)}_{G}\|\mathbb{P}_{0}^{(n)})$. ∎

### v5

@@ V Square Root Law and Approximation of the Distribution of Noise
V-B Square Root Law and Approximation of the Distribution of Noise
@@
  • With given $0<\delta<1$ and $p_{n}=c\cdot n^{-\frac{1}{2}}$ where $c$ is a constant, we can increase $n$ to satisfy any small decoding error probability $\epsilon$ without violating any proper covert constraint since the divergences between $\mathbb{P}^{(n)}_{0}$ and $\mathbb{P}^{(n)}_{1}$ will be stationary and the value depends on $c$.
@@
In this work and our previous work [23], truncated Gaussian distributions are chosen as the distributions for random coding over AWGN channels. There are two reasons of the choice. First, the way of truncation is at our choice, which will be convenient to satisfy the power constraint. The corresponding output distributions are close to Gaussian distributions with zero mean and i.i.d coordinates, hence they will lead to relatively tight normal approximations of finite blocklength throughput, which has been explained in [23]. Second, as we have explained in Section IV, truncated Gaussian distributions with proper chosen variances are good approximations of Dirac measures, which make them natural candidates on random coding for covert communications over AWGN channels. From the framework, other measures as approximations of Dirac measures can also be utilized in random coding over AWGN covert channels. Nevertheless, the number of the reliable bits from the corresponding random codes may be limited. Considering the mutual information as an approximate metric of the upper bound on the number of the transmitted bits, it is reasonable that the input distribution should be close to Gaussian.
VI Conclusion
@@
This paper presents a comprehensive analysis of covert communication over the AWGN channel, focusing on both achievability and converse bounds under finite blocklength constraints. By characterizing the first and second order asymptotics of Gaussian random coding, we establish fundamental limits on the trade-offs between reliability, covertness, and key size. Notably, the dependence of these asymptotics on the hyper-parameter $\mu$ underscores the critical role of power constraints in practical coding designs for finite $n$. Our results reveal that the required number of key bits scales as $O(\sqrt{n})$, a conclusion that remains robust even when the noise level at the legitimate receiver is marginally lower than that at the adversary. This highlights the inherent necessity of shared secrecy resources to sustain covertness in scenarios with minimal channel asymmetry. From an information-geometric perspective, the truncated Gaussian distribution emerges as the natural choice for codebook generation, balancing the trade-offs between power constraints and divergence-based covertness metrics. The Square Root Law for optimal power allocation is rigorously justified through the asymptotic behavior of divergences between Gaussian distributions and the statistical indistinguishability of noise. These findings not only unify existing theoretical frameworks but also provide practical insights for designing covert communication systems with provable guarantees.
@@ Appendix A Proof of Lemma 1
From Definition 3, an $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ is necessarily an $(M,n,\epsilon,\Psi)_{m}$ code, hence the left-hand of (16) holds. An $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ is necessarily an $(M,K,n,\epsilon,\sigma,\delta)$ code, and an $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$ is necessarily a $(M,K,n,\epsilon,\sigma,\delta)$ code, hence, we have the inequalities (17) and the right-hand of (18). The inequalities on right-hand of (16) and the left-hand of (18) follow from the fact that a code which satisfies maximal power constraint less or equal than $\Psi$ necessarily has average power less or equal than $\Psi$. ∎
Appendix B Proof of Lemma 2
Proof.
@@ Appendix B Proof of Lemma 2
$\begin{split}\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}d\bm{x}\\ =&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{1}{\Delta}d\bm{x}\\ =&-\log\Delta.\end{split}$ (163)
@@
$\mathbb{V}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=$ $\frac{1}{2}\underset{\mathbb{R}^{n}}{\int}|\bar{g}^{(n)}(\bm{x})-g^{(n)}(\bm{x})|d\bm{x}$
@@
Appendix C Proof of Lemma 3
@@
Consider the distributions $\bar{P}_{n}^{(n)}$ and $P_{n}^{(n)}$ as convolution operators $\bar{\mathfrak{F}}_{n}^{(n)}$ and $\mathfrak{F}_{n}^{(n)}$ on $C(\mathbb{R}^{n})$, respectively. We apply a multivariate version of Theorem 2 for the probability density functions $f_{0}^{(n)}(\bm{y})$ and $h_{0}^{(n)}(\bm{y})$, which are in $C_{0}(\mathbb{R}^{n})$,
$\mathfrak{F}_{n}^{(n)}f_{0}^{(n)}(\bm{y})={f}_{1}^{(n)}(\bm{y}),$ (165)
$\bar{\mathfrak{F}}_{n}^{(n)}f_{0}^{(n)}(\bm{y})=\bar{f}_{1}^{(n)}(\bm{y}),$ (166)
@@
$\mathfrak{F}_{n}^{(n)}h_{0}^{(n)}(\bm{z})=h^{(n)}_{1}(\bm{z}),$ (167)
$\bar{\mathfrak{F}}_{n}^{(n)}\bar{h}_{0}^{(n)}(\bm{z})=\bar{h}_{1}^{(n)}(\bm{z}).$ (168)
Corollary 2 and Theorem 2 imply the following statement.
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\bar{f}_{1}^{(n)}(\bm{y})-{f}_{1}^{(n)}(\bm{y})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (169)
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\bar{h}_{1}^{(n)}(\bm{z})-h^{(n)}_{1}(\bm{z})|\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty.$ (170)
As the function $\varphi(x)=\log x$ is continuous, we have
$\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}|\log\bar{f}_{1}^{(n)}(\bm{y})-\log f_{1}^{(n)}(\bm{y})|$
$=$ $\underset{\bm{y}\in\mathbb{R}^{n}}{\sup}\log\frac{\bar{f}_{1}^{(n)}(\bm{y})}{f_{1}^{(n)}(\bm{y})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (171)
@@
$\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}|\log\bar{h}_{1}^{(n)}(\bm{z})-\log h^{(n)}_{1}(\bm{z})|$
$=$ $\underset{\bm{z}\in\mathbb{R}^{n}}{\sup}\log\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\rightarrow 0\,\ \,\text{as}\,\,n\rightarrow\infty,$ (172)
@@
$\frac{\bar{f}_{1}^{(n)}(\bm{y})}{f_{1}^{(n)}(\bm{y})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty,$ (173)
$\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}\rightarrow 1\,\ \,\text{as}\,\,n\rightarrow\infty.$ (174)
@@
Appendix D Proof of Lemma 7
@@
Appendix E Proof of Lemma 10
@@
[…]
$=$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{1}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{2}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
$=$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{1}}\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{2}}d\bm{z}$
$-\int(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}})\log Q^{(n)}_{\bm{Z},0}d\bm{z}\bigg|$
$=$ $\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}$
$+\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{1}}d\bm{z}-\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log Q^{(n)}_{\bm{Z},\bm{C}_{2}}d\bm{z}$
$-\int(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}})\log Q^{(n)}_{\bm{Z},0}d\bm{z}\bigg|$
$\leq$ $\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
$+$ $\bigg|\int Q^{(n)}_{\bm{Z},\bm{C}_{2}}\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{2}}}{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}d\bm{z}\bigg|$ (176)
We analyze the two terms in (E) separately. For the first term, as the difference between $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ is the $k$-th component, we have
$\bigg|\int\big(Q^{(n)}_{\bm{Z},\bm{C}_{1}}-Q^{(n)}_{\bm{Z},\bm{C}_{2}}\big)\log\frac{Q^{(n)}_{\bm{Z},\bm{C}_{1}}}{Q^{(n)}_{\bm{Z},0}}d\bm{z}\bigg|$
@@
$\|\bm{x}\|^{2}\leq\|\bm{z}\|^{2}-\|\bm{z}-\bm{x}_{i}\|^{2}\leq\|\bm{x}\|^{2}$ (178)
and each $\bm{x}$ satisfies $\mu n\Psi(n)\leq\|\bm{x}\|^{2}\leq n\Psi(n)$.
For the second term, let
@@
then $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ can be expressed as
$Q^{(n)}_{\bm{Z},\bm{C}_{1}}=\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2},\,\ \ \ \,Q^{(n)}_{\bm{Z},\bm{C}_{2}}=\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}.$ (181)
@@
$\overset{(c)}{=}$ $\frac{\|\bm{x}_{k}-\bm{x}^{\prime}_{k}\|^{2}}{2M\sigma^{2}}$
@@
where (b) holds because KL divergence $\mathbb{D}(p\|q)$ is a convex function with the pair $(p,q)$ [38], (c) follows from the close form of KL divergence between two Gaussian distributions [39], and (d) follows from the fact both $\|\bm{x}_{k}\|^{2}$ and $\|\bm{x}^{\prime}_{k}\|^{2}$ are in $\mu^{2}n\Psi(n),n\Psi(n)]$. Combining (E), (E) and (E), we conclude
$\bigg|\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})-\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})\bigg|\leq\frac{4n\Psi(n)}{M\sigma^{2}}.$ (183)